OpenStudy (anonymous):

Solve the system. y = x + 2 and y = -x + 0

OpenStudy (anonymous):

y=-x.........(1) putting y=-x in other eqn, -x=x+2 -x-x=2 -2x=2 x=-1 as y=-x so y=-(-1)=1